# Embedding networks with the random walk first return time distribution

### v1

@@ 3.1.1 Air transit network
The world air transit network from the OpenFlights dataset 2016111Upstream: openflights.org, retrieved from networks.skewed.de/net/openflights is a network in which nodes represent airports and edges indicate the existence of at least one direct flight between them. The number and directionality of flights have been ignored, yielding an unweighted, undirected network. Airport sizes were taken from the OurAiports dataset222ourairports.com/data.
@@ 3.2 Alignment
Figure 4: Performance of three alignment algorithms on a set of 18 undirected network datasets ordered by number of edges—see Appendix C for details. For the first 15 graphs, the noisy copy is created by randomly removing $5\%$ of edges and permuting the nodes. For the last three networks, ‘real’ noise is used (see Appendix C for details). All experiments were run with the same hyperparameters ($\mu=1$) and no optimization of hyperparameters was considered.
@@ Appendix C Index of datasets for graph alignment
datasets
results
[accuracy]
[time(s)]
index name n m type FRT FUGAL FUGAL-FRT
0 ca-netscience 379 914 Collaboration
@@ results
9 socfb-Haverford76 1446 59589 Social
@@
10 socfb-Swarthmore42 1659 61050 Social
@@
11 socfb-Bowdoin47 2252 84387 Social
@@
13 socfb-Hamilton46 2314 96394 Social
@@ 17 MultiMagna* 1004 8323 Biological
Table 3: List of datasets and accuracy results used for numerical experiments on graph alignment using FRTDs and the existing FUGAL [5] algorithm previously used for a review in Ref. [38]. The indices correspond to the labels in Fig. 4. The datasets marked with an asterisk refer to ‘real noise’ networks—networks where multiple noisy copies are available and these are aligned to each other rather than to copies with artificial noise. Voles and HighSchool are temporal proximity networks, where the final network is aligned to a previous variant containing 95% of edges. MultiMagna is a protein-protein interaction network for yeast, in this case the ‘best’ network is aligned to a noisy variant containing an additional 5% of low confidence interactions.

### v2

@@ 3.1.1 Air transit network
The world air transit network from the OpenFlights dataset 2016111Upstream: openflights.org, retrieved from networks.skewed.de/net/openflights [44] is a network in which nodes represent airports and edges indicate the existence of at least one direct flight between them. The number and directionality of flights have been ignored, yielding an unweighted, undirected network. Airport sizes were taken from the OurAiports dataset222ourairports.com/data.
@@ 3.2 Alignment
Figure 4: Performance of three alignment algorithms on a set of 18 undirected network datasets ordered by number of edges—see Appendix C for details. For the first 15 graphs, the noisy copy is created by randomly removing $5\%$ of edges and permuting the nodes. For the last three networks, “real” noise is used (see Appendix C for details). All experiments were run with the same hyperparameters ($\mu=1$) and no optimization of hyperparameters was considered.
@@ Appendix C Index of datasets for graph alignment
dataset
accuracy (average $\pm$ standard deviation)
time in seconds (average $\pm$ standard deviation)
index name $\boldsymbol{n}$ $\boldsymbol{m}$ type FRT FUGAL FUGAL-FRT
0 ca-netscience [39] 379 914 collaboration
@@
9 socfb-Haverford76 [58] 1446 59589 social
@@
10 socfb-Swarthmore42 [58] 1657 61049 social
@@
11 socfb-Bowdoin47 [58] 2250 84386 social
@@
13 socfb-Hamilton46 [58] 2312 96393 social
@@ 17 Yeast* [10, 59] 1004 8323 biological
Table 3: Datasets and accuracy results for numerical experiments on graph alignment using FRTDs and the existing FUGAL [6] algorithm. We study the same set of networks previously used in the review of Ref. [53]. Some of these versions of the networks are different to the original data, e.g., they are slightly different sizes. We could not tell exactly how they are different but use the same versions as Ref. [53] for consistency (retrieved from github.com/constantinosskitsas/Framework_GraphAlignment). Data are also available in the repositories of Refs. [49, 44, 35]. Indices correspond to the labels in Fig. 4. Networks marked with an asterisk correspond to “real noise” networks—networks where multiple noisy copies are available and these are aligned to each other rather than to copies with artificial noise. Voles and HighSchool are temporal proximity networks, where the final network is aligned to a previous variant containing 95% of edges (we acknowledge the SocioPatterns collaboration for the HighSchool network). Yeast is a protein-protein interaction network for yeast, in this case the “best” network is aligned to a noisy variant containing an additional 5% of low confidence interactions.